OpenStudy (anonymous):

Jeff stores coffee powder in a steel container with a capacity of 252.3 cubic inches. He uses a glass container to store sugar. The two containers are congruent cylinders. If the glass container has a height of 6.2 inches, what is its radius? Answer 1.8 inches 2.1 inches 3.1 inches 3.6 inches

OpenStudy (campbell_st):

the volume of a cylinder is \[V = \pi r^2 h\] you know V and h substitute and find r \[252.3 = \pi r^2 \times 6.2\] \[r^2 = \frac{252.3}{6.2\pi}\] I'll let you finish off the question.

OpenStudy (anonymous):

no no no you don't have to do that you can finish it lol jk but really i worked it out and i think i did something wrong my answer didnt come up as one of the options...